Students arrive for peer tutoring lab an average of 4.3 students per hour. What is the probablitly that less than 2 students will show up in an hour?
A) 0.8385
B) 0.0719
C) 0.9281
D) 0.1615
Answer #1

Ans:

Use Poisson distribution with mean rate=4.3 per hr

P(x=k)=e-4.3*(4.3^k/k!)

P(x<2)=P(x=0)+P(x=1)

=e-4.3*(4.3^0/0!+4.3^1/1!)

=e-4.3*(1+4.3)

=0.0719

Option B is correct.(0.0719)